OpenStudy (anonymous):

In rectangle PQRS, PR = 18x - 28and QS = x + 380. Find the value of x and the length of each diagonal.

OpenStudy (ranga):

The two diagonals of a rectangle are of the same length. So you can equate PR and QS and solve for x.

OpenStudy (anonymous):

But how to solve for X exactly?

OpenStudy (mertsj):

Write the equation as ranga instructed. What is it?

OpenStudy (anonymous):

in a rectangle both diagonals are equal, so you need to equate both the diagonals ie: 18x - 28 = x + 380 or 17x = 408 = 24 no each diagonal would be 24+380 = 404

OpenStudy (anonymous):

THANKYOU!

OpenStudy (anonymous):

ur welcome